OpenStudy (anonymous):

What is the answer for 10= x/4 - 8

OpenStudy (anonymous):

"no answers" but it is easy work backwards. 8 PLUS ten equals 18. Now 18 TIMES 4 equals what? that's the answer

OpenStudy (x3_drummerchick):

is the problem  ?

OpenStudy (anonymous):

72 10=x/4-8 +8 +8 -------------- 18=x/4 18(*4)=(x/4)*(4) -------------- 72=x

OpenStudy (anonymous):

or just tell her the answer okay thanks

OpenStudy (anonymous):

im with drummerkick, if the problem is 10=(x)/(4-8), the answer will be different

OpenStudy (x3_drummerchick):